# Can Generative Geospatial Diffusion Models Excel as Discriminative Geospatial Foundation Models?

### v1

@@ Abstract
Self-supervised learning (SSL) has revolutionized representation learning in Remote Sensing (RS), advancing Geospatial Foundation Models (GFMs) to leverage vast unlabeled satellite imagery for diverse downstream tasks. Currently, GFMs primarily focus on discriminative objectives, such as contrastive learning or masked image modeling, owing to their proven success in learning transferable representations. However, generative diffusion models—which demonstrate the potential to capture multi-grained semantics essential for RS tasks during image generation—remain underexplored for discriminative applications. This prompts the question: can generative diffusion models also excel and serve as GFMs with sufficient discriminative power? In this work, we answer this question with SatDiFuser, a framework that transforms a diffusion-based generative geospatial foundation model into a powerful pretraining tool for discriminative RS. By systematically analyzing multi-stage, noise-dependent diffusion features, we develop three fusion strategies to effectively leverage these diverse representations. Extensive experiments on remote sensing benchmarks show that SatDiFuser outperforms state-of-the-art GFMs, achieving gains of up to +5.7% mIoU in semantic segmentation and +7.9% F1-score in classification, demonstrating the capacity of diffusion-based generative foundation models to rival or exceed discriminative GFMs. Code will be released.
@@ 2 Related Work
A number of recent works [64, 60, 38, 31, 50, 9, 22, 67, 11] in computer vision have started investigating the discriminative representations inherent in pretrained diffusion models. Some methods mine self-attention tensors for unsupervised segmentation, either by merging attention maps (DiffSeg [50]) or by constructing affinity graphs (DiffCut [9]). Others incorporate cross-attention signals (SLiMe [22]), fine-tuning text embeddings to segment objects at varied granularity. Meanwhile, DatasetDM [60] extracts multi-scale features from a Stable Diffusion UNet to train a dataset-generation model capable of producing densely annotated images. Diffusion HyperFeatures [31] further enhances feature aggregation by incorporating multi-timestep feature maps, creating a feature descriptor for semantic keypoint correspondence tasks. Additionally, REPA [65] demonstrates the improving synergies between representation learning and generative models by utilizing external high-quality representations. Inspired by these efforts, we exploit diffusion models that are trained on large-scale global satellite imagery, adapting their representational capacity to a wide range of RS tasks.
@@
Diffusion models have gained increased traction in RS, being applied to image generation, enhancement, and interpretation [29]. Recent efforts [23, 66, 69, 48] have focused on developing diffusion-based generative foundation models for high-fidelity satellite image synthesis. For instance, DiffusionSat [23] generates data conditioned on semantic text and metadata, while MetaEarth [66] enables arbitrary-sized image generation using a resolution-guided approach. Beyond synthesis, numerous diffusion-based methods address image enhancement tasks, including denoising [17, 39], cloud removal [56, 71], and super-resolution [54, 10], showcasing their versatility in RS. Another line of research focuses on discriminative applications [26], though these often rely on labeled data and are limited to specific tasks, such as semantic segmentation [1, 24, 70, 27, 40] or change detection [59, 51, 68, 21]. For example, SegDiff [1] diffuse ground-truth masks, while others use class predictions or labeled guidance [24, 27, 40]. Although a few studies have explored diffusion as a label-free pretraining framework, they remain narrowly focused on a single application scenario, such as hyperspectral images segmentation [70], or change detection [4]. In contrast, our work provides a comprehensive investigation of the discriminative capabilities of diffusion-based generative models across multiple RS tasks. By moving beyond task-specific solutions and limited testing, our method advances the broader potential of diffusion-driven GFMs pretrained on global-scale data.
@@ 3.1 Feature Extraction from Diffusion Process
In an LDM [43], an input image $\mathbf{x}$ is first mapped into a latent representation $\mathbf{z}\in\mathbb{R}^{H_{0}\times W_{0}\times C_{0}}$ via an autoencoder. To extract features from the diffusion process, we start with the clean latents $\mathbf{z}$ and employ DDIM inversion [46] to trace a reverse noise path, obtaining noisy latents. We then run the denoising diffusion model on these noisy latents to capture multi-scale multi-timestep feature maps. This inversion approach yields faithful latent representations, helping preserve fine-grained details in deterministic tasks. Complete equations are provided in the supplementary material.
@@ 4.1 Evaluation Protocol
We adopt GEO-Bench [25], which contains semantic segmentation and classification datasets, covering diverse application domains (e.g., agriculture, urban, forest, etc.) and geographic regions. Dataset-specific details, including dataset sizes and spectral properties, are summarized in the supplementary material.
@@ 4.2 Implementation Details
For DiffusionSat, we select ResNet and self-attention outputs at timesteps $\{1,100,200\}$ from the decoder blocks of its UNet (details justified in Sec. 4.4). For comparison GFMs, we extract features from evenly spaced layers based on the specific GFM architecture, following common protocols [35, 42]. For instance, in a 12-layer ViT-based MAE model, we select features from layers indexed at (3, 5, 7, 11). All models are optimized using AdamW with an initial learning rate of 0.01, scheduled with cosine decay after a 5-epoch warm-up. Images are cropped or resized to match the pretraining resolution required by each GFM; specifically for DiffusionSat, which offers models trained on 512px and 256px resolutions, we resize images to these dimensions using bilinear interpolation based on their original resolution. Each spectral band is normalized individually using the minimum and maximum values calculated across the entire dataset. We use a training batch size of 32. Additional dataset-specific implementation details, such as loss functions and training epochs, are provided in the supplementary material.
@@ 4.4 Ablation Studies
Using one classification and one segmentation dataset as examples (Fig. 3), we observe that performance peaks when sampling within the first  20% of the diffusion timesteps. Traditionally, DDIM [46] sampling can take up to 1000 steps. At later timesteps, heavily noised latents lose too many fine-grained details, while very early steps (e.g., one-step denoising) provide limited learning signals.
@@
Table 3: Comparion of using raw features and different feature fusion strategies. The reported numbers are top-1 accuracy or mIoU.
@@
We quantify the impact on the number of experts $E$ and routing parameter top-$k$ in MoE fusion method. Tab. 4 shows that fewer experts lead to insufficient learning, while more experts may introduce redundancy. Using top-$k$=2 consistently outperforms a single expert activation, suggesting the benefit of complementary expert representations. We select $E=8$ and top-$k$=2 as our configuration to balance computational efficiency and performance.

### v2

@@ Abstract
Self-supervised learning (SSL) has revolutionized representation learning in Remote Sensing (RS), advancing Geospatial Foundation Models (GFMs) to leverage vast unlabeled satellite imagery for diverse downstream tasks. Currently, GFMs primarily employ objectives like contrastive learning or masked image modeling, owing to their proven success in learning transferable representations. However, generative diffusion models, which demonstrate the potential to capture multi-grained semantics essential for RS tasks during image generation, remain underexplored for discriminative applications. This prompts the question: can generative diffusion models also excel and serve as GFMs with sufficient discriminative power? In this work, we answer this question with SatDiFuser, a framework that transforms a diffusion-based generative geospatial foundation model into a powerful pretraining tool for discriminative RS. By systematically analyzing multi-stage, noise-dependent diffusion features, we develop three fusion strategies to effectively leverage these diverse representations. Extensive experiments on remote sensing benchmarks show that SatDiFuser outperforms state-of-the-art GFMs, achieving gains of up to +5.7% mIoU in semantic segmentation and +7.9% F1-score in classification, demonstrating the capacity of diffusion-based generative foundation models to rival or exceed discriminative GFMs. The source code is available at: https://github.com/yurujaja/SatDiFuser.
@@ 2 Related Work
A number of recent works [5, 69, 65, 64, 40, 33, 53, 10, 24, 72, 13] in computer vision have started investigating the discriminative representations inherent in pretrained diffusion models. Some methods mine self-attention tensors for unsupervised segmentation, either by merging attention maps (DiffSeg [53]) or by constructing affinity graphs (DiffCut [10]). Others incorporate cross-attention signals (SLiMe [24]), fine-tuning text embeddings to segment objects at varied granularity. Meanwhile, DatasetDM [64] extracts multi-scale features from a Stable Diffusion UNet to train a dataset-generation model capable of producing densely annotated images. Diffusion HyperFeatures [33] further enhances feature aggregation by incorporating multi-timestep feature maps, creating a feature descriptor for semantic keypoint correspondence tasks. Additionally, REPA [70] demonstrates the improving synergies between representation learning and generative models by utilizing external high-quality representations. Inspired by these efforts, we exploit diffusion models that are trained on large-scale global satellite imagery, adapting their representational capacity to a wide range of RS tasks.
@@
Diffusion models have gained increased traction in RS, being applied to image generation, enhancement, and interpretation [31]. Recent efforts [25, 71, 74, 51, 55, 12] have focused on developing diffusion-based generative foundation models for high-fidelity satellite image synthesis. For instance, DiffusionSat [25] generates data conditioned on semantic text and metadata, while MetaEarth [71] enables arbitrary-sized image generation using a resolution-guided approach. Beyond synthesis, numerous diffusion-based methods address image enhancement tasks, including denoising [19, 41], cloud removal [60, 76], and super-resolution [58, 11], showcasing their versatility in RS. Another line of research focuses on discriminative applications [28], though these often rely on labeled data and are limited to specific tasks, such as semantic segmentation [1, 26, 75, 29, 42] or change detection [63, 54, 73, 23]. For example, SegDiff [1] diffuse ground-truth masks, while others use class predictions or labeled guidance [26, 29, 42]. Although a few studies have explored diffusion as a label-free pretraining framework, they remain narrowly focused on a single application scenario, such as hyperspectral images segmentation [75, 48], or change detection [4]. In contrast, our work provides a comprehensive investigation of the discriminative capabilities of diffusion-based generative models across multiple RS tasks. By moving beyond task-specific solutions and limited testing, our method advances the broader potential of diffusion-driven GFMs pretrained on global-scale data.
@@ 3.1 Feature Extraction from Diffusion Process
In an LDM [45], an input image $\mathbf{x}$ is first mapped into a latent representation $\mathbf{z}\in\mathbb{R}^{H_{0}\times W_{0}\times C_{0}}$ via an autoencoder. To extract features from the diffusion process, we start with the clean latents $\mathbf{z}$ and employ DDIM inversion [49] to trace a reverse noise path, obtaining noisy latents. We then run the denoising diffusion model on these noisy latents to capture multi-scale multi-timestep feature maps. This inversion approach yields faithful latent representations, helping preserve fine-grained details in deterministic tasks. Complete equations are provided in the Supplementary Sec. 6.
@@ 4.1 Evaluation Protocol
We adopt GEO-Bench [27], which contains semantic segmentation and classification datasets, covering diverse application domains (e.g., agriculture, urban, forest, etc.) and geographic regions. Dataset-specific details, including dataset sizes and spectral properties, are summarized in the supplementary material (Sec. 7).
@@ 4.2 Implementation Details
For DiffusionSat, we select ResNet and self-attention outputs at timesteps $\{1,100,200\}$ from the decoder blocks of its UNet (details justified in Sec. 4.4). For comparison GFMs, we extract features from evenly spaced layers based on the specific GFM architecture, following common protocols [37, 44]. For instance, in a 12-layer ViT-based MAE model, we select features from layers indexed at (3, 5, 7, 11). All models are optimized using AdamW with an initial learning rate of 0.01, scheduled with cosine decay after a 5-epoch warm-up. Images are cropped or resized to match the pretraining resolution required by each GFM; specifically for DiffusionSat, which offers models trained on 512px and 256px resolutions, we resize images to these dimensions using bilinear interpolation based on their original resolution. Each spectral band is normalized individually using the minimum and maximum values calculated across the entire dataset. We use a training batch size of 32. Additional dataset-specific implementation details, such as loss functions and training epochs, are provided in the supplementary material (Sec. 8).
@@ 4.4 Ablation Studies
Using one classification and one segmentation task as examples (Fig. 3), we observe that performance peaks when sampling within the first  20% of the diffusion timesteps. Typically, DDIM [49] sampling can take up to 1000 steps. At later steps, heavily noised latents lose too many fine-grained details, while very early steps (e.g., one-step denoising) provide limited learning signals.
@@
Table 3: Comparison of using raw features and different feature fusion strategies. The reported numbers are top-1 accuracy or mIoU.
@@
We quantify the impact on the number of experts $E$ and routing parameter top-$k$ in MoE fusion method. Tab. 4 shows that fewer experts lead to insufficient learning, while more experts may introduce redundancy. Using top-kk=2 consistently outperforms a single expert activation, suggesting the benefit of complementary expert representations. We select $E=8$ and top-kk=2 as our configuration to balance computational efficiency and performance.
@@ 5 Conclusion
Acknowledgements
This project is funded by the Division of Geoinformatics at KTH Royal Institute of Technology, the Geomatics section at KU Leuven and the Natural Science Foundation of Shanghai (Grant No. 25ZR1402268). We also thank the National Academic Infrastructure for Supercomputing in Sweden (NAISS, Grant No. 2022-06725) for supporting the computations and data handling.
Supplementary Material
6 Diffusion Models
Below, we provide a concise mathematical overview of discrete diffusion models (DMs).
6.1 Diffusion Model Essentials
Let $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$ be a sample from an underlying data distribution. A forward diffusion process iteratively adds Gaussian noise over $T$ discrete timesteps, producing corrupted samples $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$. One common choice to model each step is:
$p(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}\!\bigl{(}\mathbf{x}_{t};\,\sqrt{1-\beta_{t}}\,\mathbf{x}_{t-1},\,\beta_{t}\,\mathbf{I}\bigr{)},$ (5)
where $\beta_{t}\in(0,1)$ controls the noise variance. One can obtain $\mathbf{x}_{t}$ directly from the original image $\mathbf{x}_{0}$ given the cumulative effect of t nose-adding steps:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\,\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon},\quad\bar{\alpha}_{t}=\prod_{i=1}^{t}{\alpha_{i}},$ (6)
where $\alpha_{t}=1-\beta_{t}$ and $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. As $t$ grows, $\mathbf{x}_{t}$ becomes increasingly noisy; at $t=T$, the corrupted $\mathbf{x}_{T}$ approximates a pure Gaussian distribution, losing most structure of $\mathbf{x}_{0}$.
To generate novel samples starting from pure noise, a diffusion model learns a reverse denoising process $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$, parameterized by $\theta$, which conceptually “denoises” $\mathbf{x}_{t}$ step by step until recovering $\mathbf{x}_{0}$:
$p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}\!\bigl{(}\mathbf{x}_{t-1};\,\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t),\,\Sigma_{\theta}(\mathbf{x}_{t},t)\bigr{)}.$ (7)
A common training objective is to minimize the distance between the true noise $\boldsymbol{\epsilon}_{t}$ and the model’s predicted noise $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$:
$\mathcal{L}_{\text{simple}}(\theta)\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\epsilon},t}\bigl{[}\|\boldsymbol{\epsilon}_{t}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\|^{2}\bigr{]}.$ (8)
At inference, sampling proceeds from $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and iteratively applies $p_{\theta}$ to yield a final $\mathbf{x}_{0}$.
6.2 Feature Extraction using Diffusion Models
While diffusion models (DMs) are primarily designed for image generation from Gaussian noise, our goal is to extract their learned representations for real images. To achieve this, we first invert a real image into a noisy state and then perform the reverse denoising process.
To illustrate the inversion step, we revisit DDIM [49], a widely adopted sampling approach known for faster generation and invertibility. A common deterministic formulation for going from $\mathbf{x}_{t}$ to $\mathbf{x}_{t-1}$ is:
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\hat{\mathbf{x}_{0}}+\sqrt{1-\alpha_{t-1}}\;\epsilon_{\theta}(\mathbf{x}_{t},t).$ (9)
where $\hat{\mathbf{x}_{0}}=\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\,\epsilon_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\alpha_{t}}}$ is the predicted clean image. By removing explicit Gaussian noise additions at each step, the process becomes deterministic, allowing a “mirror pass” that encodes $\mathbf{x}_{0}$ to $\mathbf{x}_{T}$. If we then use $\mathbf{x}_{T}$ as the start of the usual sampling procedure, we recover the original $\mathbf{x}_{0}$.
We leverage this property by reversing the order of timesteps, going from $\mathbf{x}_{t-1}$ to $\mathbf{x}_{t}$, starting from $\mathbf{x}_{0}$, and then running the denoising process on $\mathbf{x}_{t}$ to extract features.
While one could alternatively introduce noise into a real image by selecting a timestep and manually adding noise via Eq. 6, this approach introduces stochastic variations. To ensure consistency, we adopt DDIM inversion, leveraging its deterministic nature to repurpose diffusion models for discriminative tasks.
7 Datasets Description
There are six classification datasets in GEO-Bench [27]:
m-bigearthnet It contains 120 $\times$ 120 images with 43 land cover classes. The dataset includes 20,000 training samples, 1,000 validation samples, and 1,000 test samples. It consists of 12 spectral bands obtained from Sentinel-2 imagery, with a spatial resolution of 10.0m for the RGB channels.
m-brick-kiln It consists of 64 $\times$ 64 images with 2 classes, focusing on brick kiln detection. The dataset includes 15,063 training samples, 999 validation samples, and 999 test samples. The imagery is derived from Sentinel-2 with 10 spectral bands and a resolution of 10.0m for RGB. Additional Sentinel-1 data is included.
m-eurosat It consists of 64 $\times$ 64 images spanning 10 land cover classes. The dataset contains 2,000 training samples, 1,000 validation samples, and 1,000 test samples. It includes 13 spectral bands captured from Sentinel-2 with an RGB resolution of 10.0m.
m-forestnet This dataset contains 332 $\times$ 332 images and covers 12 classes related to forest monitoring. It includes 6,464 training samples, 989 validation samples, and 993 test samples. The dataset comprises 6 spectral bands obtained from Landsat-8, with a spatial resolution of 15.0m for the RGB channels.
m-pv4ger This dataset comprises 320 $\times$ 320 images covering 2 classes, with 11,814 training samples, 999 validation samples, and 999 test samples. The imagery is obtained from RGB data, with a spatial resolution of 0.1m.
m-so2sat This dataset consists of 32 $\times$ 32 images spanning 17 different land cover classes. It contains 19,992 training samples, 986 validation samples, and 986 test samples. The images are derived from Sentinel-2 data with 13 spectral bands and a spatial resolution of 10.0m for RGB.
In addition, six semantic segmentation datasets are included:
m-pv4ger-seg It is the segmentation variant of m-pv4ger, containing 320 $\times$ 320 images with 3,000 training samples, 403 validation samples, and 403 test samples. The dataset has 3 spectral bands (RGB) with a spatial resolution of 0.1m.
m-nz-cattle It contains 500 $\times$ 500 images with 2 classes, including 524 training samples, 66 validation samples, and 65 test samples. The imagery consists of 3 spectral bands (RGB) with an unknown spatial resolution.
m-NeonTree It includes 400 $\times$ 400 images with 2 classes, consisting of 270 training samples, 94 validation samples, and 93 test samples. The dataset comprises 5 spectral bands (RGB + Hyperspectral + Elevation).
m-cashew-plantation It comprises 256 $\times$ 256 images with 7 classes, featuring 1,350 training samples, 400 validation samples, and 400 test samples. The imagery is sourced from Sentinel-2 with 10 spectral bands and an RGB resolution of 10.0m.
m-SA-crop-type This dataset consists of 256 $\times$ 256 images with 10 classes. It contains 3,000 training samples, 1,000 validation samples, and 1,000 test samples. The imagery is sourced from Sentinel-2 with 10 spectral bands and an RGB resolution of 10.0m.
m-chesapeake-landcover This dataset consists of 256 $\times$ 256 images with 7 land cover classes. It contains 3,000 training samples, 1,000 validation samples, and 1,000 test samples. The dataset includes 4 spectral bands (RGBN) with a spatial resolution of 1.0m.
8 Evaluation Details
We provide additional details regarding the evaluation process in this section
For task training criteria, we utilize nn.CrossEntropyLoss() from the PyTorch library for all tasks, except for the m-bigearthnet dataset, which follows a multi-label classification setup and nn.BCEWithLogitsLoss() is applied.
Regarding training schedules, all models are trained for 60 epochs on the m-cashew-plantation and m-sa-crop-type datasets, while the remaining datasets undergo training for 40 epochs.
Additionally, we apply data augmentation techniques for all datasets to enhance model generalization. During training, images undergo random horizontal flipping, vertical flipping, and color jittering based on a probabilistic threshold of 0.5.
9 Visualizations of Global Weighting
(a) Weight allocations for m-chesapeake-landcover dataset.
(b) Weight allocations for m-pv4ger-seg dataset.
(c) Weight allocations for m-NeonTree dataset.
Figure 7: Normalized weight allocations in Global Weighted Fusion across different datasets.
In Sec. 4.4, we demonstrate that features from different blocks and timesteps contribute differently depending on the dataset. Our global weighted fusion method effectively aggregates these features to enhance performance. Fig. 7 visualizes the learned weight distributions across blocks and timesteps at different scales, illustrating how our fusion strategy dynamically adjusts feature importance for each dataset. This automated weighting helps reduce the need for manual feature selection, promoting adaptive and optimal feature integration.